QVX, Inc. manufactures an intermediate part it uses for assembling its finished

product. QCX normally produces 2,500 units of the product. Manufacturing costs are:

direct materials $1,20; direct labor $2.00; variable overhead $1.50; fixed overhead

allocated $2.50 per unit. An outside supplier offered to sell a similar part at $5.00 per

unit QVX will leave the facilities idle even if the part is purchased from outside.

a. Should the company make or buy the part?

b. The facility that is released when QVX does not produce the part can be

rented for $1,000 for the period. Should the company make or buy the part?
Answer #1

a. When the company has the choice of making a product or buying it, only the relevant costs are considered while making calculations because unavoidable or allocated overhead cost will be still there. .Relevant costs are the costs that are incurred only when making specific business decisions.

In short, the additional costs for making an additional unit is the relevant cost. All the variable costs are the relevant costs as they are incurred on the basis of units produced. If the specific fixed cost in incurred for the new project, it also becomes the relevant cost as it is spent only for that new project. All the other fixed costs which were spent even before taking up the new projects are the irrelevant costs.

Therefore, the relevant cost of production of 2,500 units:

Particulars Amount
Direct Material per unit A $1.20
Direct Labor per unit B $2.00
Variable overhead per unit C $1.50
Relevant cost per unit D=A+B+C $4.70
Total units E 2,500
Relevant cost 2,500 units F =D*E $11,750
Buying cost for 2,500 units G = 2,500 units * $5 $12,500

As the relevant cost of production i.e. $11,750 is less than the purchasing cost i.e. $12,500, hence the company should make the product.

b. Overall benefit in making the product = $12,500 - $11,750 = $750

Now if the company can rent the place for $1,000, then the position will change as the company will definitely gain $750 by producing it but it will not able to generate the rent of $1,000 and will be termed as opportunity cost.

Therefore, overall benefit of buying= Opportunity cost + Benefit/(loss) in making the product

Therefore, overall benefit of buying = $1,000 + ($750) = $250.

Hence, the company should now but the product.